# AIDE: Task-Specific Fine Tuning with Attribute Guided MultI-Hop Data Expansion

### v1

@@ Appendix A Benchmark Statistics
Appendix B Prompt for Extracting A Topic and Knowledge Attributes
@@
Appendix C Prompt for Synthesizing Data Points with A Triplet and An Operation
@@
Appendix D Prompt for Synthesizing Data Points with A Topic and Personas
@@
Appendix E An Example of 10-hop Synthesis without Residual Connection
@@
Appendix F An Example of 10-hop Synthesis with Residual Connection
@@
Appendix G Prompt for Self-reflection
@@
Appendix H Ethical Considerations
@@ Appendix I Limitations
Cognitive Process. While AIDE helps base models improve their performance in the Math task, the zero-shot performance of the fine-tuned base models remain around 25%. In the future, a potential future direction is to integrate Chain-of-Thought techniques into AIDE, such that AIDE can provide better synthetic data to enhance reasoning steps of the base models though fine-tuning.

### v2

@@ Appendix H Benchmark Statistics
  • MedQA Jin et al. (2020) is a comprehensive resource designed to enhance medical question-answering systems. It comprises 10,178 multiple-choice questions sourced from medical exams across the United States, Mainland China, and Taiwan. Each question is accompanied by several answer options, with the correct answer clearly indicated. We select 1,956 data points for the training set and 217 for the validation set. Additionally, we sample 10 seed data points to synthesize 2,173 data points through AIDE.
  • FinBen Xie et al. (2024) is part of the PIXIU project Xie et al. (2023), an open-source initiative aimed at developing, fine-tuning, and evaluating large language models (LLMs) in the financial domain. PIXIU encompasses various components including FinBen, a financial language benchmark. The CFA task consists of 1.03k data points, which we divide as follows: 100 data points for the test set, 804 as gold training data, 89 for the validation set, and 10 as seed data points to synthesize 893 additional data points through AIDE.
Task Name Abbreviation # Test data
Code Line Descriptions Code 60
Cause and Effect C&E 153
Implicatures Impl. 492
Elementary Math Math 7,688
Temporal Sequence Time 1,000
High School Biology Bio. 300
College Computer Science CS 100
Philosophy Phi. 311
Electrical Engineering EE 145
Marketing Market. 234
Flare-cfa CFA 100
ARC-Challenge - 1,170
GSM8k - 1,320
TruthfulQA - 817
MedQA - 100
Table 8: Data statistic of selected tasks from BIG-Bench, MMLU, ARC-Challenge, GSM8K and Truthful QA.
Appendix I Prompt for Extracting a Topic and Knowledge Attributes
@@
Appendix J Prompt for Synthesizing Data Points with a Triplet and an Operation
@@
Appendix K Prompt for Synthesizing Data Points with a Topic and Personas
@@
Appendix L An Example of 10-hop Synthesis without Residual Connection
@@
Appendix M An Example of 10-hop Synthesis with Residual Connection
@@
Appendix N Prompt for Self-Reflection
@@
Appendix O Ethical Considerations
@@
Cognitive Process. While AIDE helps base models improve their performance in the Math task, the zero-shot performance of the fine-tuned base models remain around 20%. In the future, a potential future direction is to integrate Chain-of-Thought techniques into AIDE, such that AIDE can provide better synthetic data to enhance reasoning steps of the base models though fine-tuning.